How many grams of 60% silver should be added to 2 grams of 20% silver to produce a 35% silver alloy?

so i tried to make a table
sliver amount % total quantity
silver 1 x 60% .6x
silver 2

Answer by checkley77(12844)

.60X+.20*2=.35(X+2)
.60X+.40=.35X+.70
.60X-35X=.7-.4
.25X=.3
X=.3/.25
X=1.2 G. OF 60% SILVER IS USED.
PROOF:
.60*1.2+.2*2=.35(1.2+2)
.72+.4=.35*3.2
1.12=1.12